OpenStudy (anonymous):

Solve x2 - 6x = 27 by completing the square. Show your work for full credit.

OpenStudy (anonymous):

(x-3)(x+9)

jimthompson5910 (jim_thompson5910):

Hint: take half of -6 to get -3 square -3 to get 9 add this to both sides

jimthompson5910 (jim_thompson5910):

x^2 - 6x = 27 x^2 - 6x + 9 = 27 + 9 x^2 - 6x + 9 = 36 I'll let you finish